# The CCube reconstruction algorithm for the SoLid experiment

### v1

@@ Abstract
The SoLid experiment is a very-short-baseline experiment aimed at searching for nuclear reactor-produced active to sterile antineutrino oscillations. The detection principle is based on the pairing of two types of solid scintillators: polyvinyl toluene and 6LiF:ZnS(Ag), which is a new technology used in this field of Physics. In addition to good neutron-gamma discrimination, this setup allows the detector to be highly segmented (the basic detection unit is a 5 cm side cube). High segmentation provides numerous advantages, including the precise location of Inverse Beta Decay (IBD) products, the derivation of the considerate antineutrino energy estimator, and a powerful background reduction tool based on the topological signature of the signal. Finally, the system is read out by a network of wavelength-shifting fibres coupled to a photodetector (MPPC). This paper describes the design of the reconstruction algorithm that allows maximum use of the granularity of the detector. The goal of the algorithm is to convert the output of the optical-fibre readout to the list of the detection units from which it originated. This paper provides a performance comparison for three methods and concludes with a choice of the baseline approach for the experiment.
@@
0000 , 1111MSC:
0000 , 1111 2023 by SoLid collaboration is licensed under CC BY 4.0[inst1] organization=Université Clermont Auvergne, CNRS/IN2P3, LPCA, city=Clermont-Ferrand, country=France
@@ 1 Introduction
The SoLid experiment is located in the vicinity of the BR2 research reactor at the SCK CEN site in Mol, Belgium. The experiment aims to make a precise measurement of the reactor antineutrino flux at a very short baseline (6.7 - 9.2 m), as well as its energy spectrum, to study the “5 MeV bump” [1]. The first measurement allows, in principle, information on the so-called reactor anomaly RAA [2] and the Gallium anomaly [3] to be provided. In particular, it is possible to test the 3+1 model [4], which assumes the existence of an additional light-sterile neutrinos state. The SoLid experiment was designed to probe the best-fit region of the oscillation parameters with sin() 0.1 and m = 1 eV2.
@@
The description of the design of the SoLid detector is beyond the scope of this article. The interested reader can consult the detailed discussion reported in [5]. The following paragraph contains the executive summary required for the description of the reconstruction algorithm.
@@
The basic detection unit of the detector is a 5 cm side polyvinyl toluene (PVT) cube. The PVT is a plastic scintillator that acts as a proton-enriched target for the antineutrino and detects the light generated by the IBD positron and consequent annihilation gamma. The detection units are combined into planes of 1616 units each. Each cube is individually wrapped in Tyvek to prevent scintillation light from escaping. Furthermore, each plane is surrounded by two square Tyvek sheets, which optically decouple the planes by preventing light from passing between them even further (shown in Figure 1). The latter is a very important feature of the calibration algorithm. Ten planes make up a module, and there are five modules in total. The scintillation light from the PVT cube is collected by two vertical and two horizontal wavelength-shifting (WLS) fibres that pass through each detection cell. One side of each WLS fibre is covered with a Mylar foil that acts as a mirror to reflect the incoming light towards the other side. The second side is coupled with the multi-pixel photon counters (MPPCs) that read out the light. The digitised version of the readout received from the detector is the starting point for any further analysis. Therefore, signal candidates must be defined from this input. The most convenient and useful way to represent the data is to transform the MPPC response back into the list of detection cubes involved in the event. This type of problem, i.e. image reconstruction, has been investigated in medical physics, for example, in [6].
@@
Several algorithms have been applied for the reconstruction of SoLid data. Two of them are already well-known and widely used. The first is a regularisation approach called the Fast Iterative Shrinkage Threshold Algorithm or FISTA, which is an improved version of ISTA [7]. The second is the Bayesian Maximum-Likelihood Expectation-Minimisation (ML-EM) approach, which is used e.g. in the NEXT experiment [8]. Finally, the last method is custom-made and builds on the ML-EM algorithm by augmenting it with information from the underlying physics processes at work. All three of them will be presented in this article, and their performances will be compared.
@@
One last important thing to mention is that all the results presented in this paper are obtained solely with Monte Carlo (MC) simulations. The production of simulation files in the SoLid experiment is based on several components. First, a detailed simulation of the core of the BR2 reactor is produced. It couples the reference ILL spectrum, fission rate predictions extracted from the MCNPX 3D model of the reactor core combined with the CINDER90 evolution code, and the MURE code to track the burn-up of the fissile products. Second, a detailed Geant4 [9] simulation of the geometry of the SoLid detector is performed. Third, the SoLid Oscillation (SoLO) framework, which works as an antineutrino generator and mediator between the two simulations introduced above. For each signal event, SoLO peaks a position inside the reactor core according to the given fission map, an antineutrino energy from the predicted spectrum, and the interaction point inside the detector. By construction, this approach takes into account the geometrical acceptance of the detector. Thus, the computational cost is significantly reduced compared to isotropic generation. To reduce it even further, SoLO directly transfers the kinematic information of the positron and neutron to the detector simulation. The interested reader can consult the detailed description of the SoLO framework reported in [10]. The final piece that handles the output of the Geant4 simulation is the so-called ReadOut Simulation (ROSim). The goal of ROSim is to reproduce the response from the detector. Practically, it transforms the energy deposits into the number of scintillation photons that are read out by the correspondent MPPCs. The detailed tuning of the ROSim parameters is reported in [11].
@@ 2 Signal definition
The SoLid detector uses the IBD process to detect antineutrino. This justifies the use of two scintillators: 6LiF:ZnS(Ag) screens to capture the neutron and PVT as a proton-enriched target for the antineutrino, hence recording the positron scintillation. For the former, once the neutron is thermalised through elastic collision, it is captured and produces tritium and particles, detected in the ZnS(Ag) crystal, which is integrated in the neutron detection screens. The latter detects in addition to the positron scintillation the subsequent positron annihilation gammas. The antineutrino energy is defined from the IBD process as follows:
@@ 3 The CCube algorithm
Let us mathematically define the problem. By postulating aij as the projector from cube j to MPPC i, the readout value is calculated as follows:
@@
Equation 3 is a discrete linear inverse problem. The problem is well-posed if it meets certain requirements: existence, uniqueness, and stability of the solution with respect to the small shifts. In the current case, the uniqueness condition is violated, as illustrated in Figure 3. For simplicity, each cube communicates there only through the fibres that cross it: the projectors aij of the cubes crossed by the given fibre i are set to 1, and to 0 for all other pairs of i,j. This indicates that only two energy deposits in the same plane are enough to allow the readout to be fulfilled in several ways. In particular, false cubes may be created at the intersection of the fibres of the true ones. In addition to the multiplicity of events, false cubes also introduce an energy bias to the initial energy deposits. Thus, the reconstruction problem is ill-posed and shall be solved either by regularisation or Bayesian methods. Before discussing the methods themselves, let us define the set of estimators that are used to assess the performance of different methods.
@@ 4 Reconstruction procedure estimators
The estimators used to evaluate the performance of the various methods are divided into two categories. The first is related to the convergence of the algorithms only. The estimators in this group calculate the deviation (also indicated as data fidelity) of the solution derived (E) from the actual measurements (p). The least squares norm and the Kullback-Leiber divergence [12] were selected to obtain it. Their mathematical definition reads as follows:
@@
The second group of estimators involves physical performance. Geant4 (G4) simulation serves as truth-level information, while the output of the SoLid Analysis Framework (Saffron2 [13] or simply S2) with the reconstruction approach at work serves as a derived solution. The comparison relies on the list of cubes per event, their coordinates (positions in the plane) and associated energy contributions. To begin with, the total number of cubes in which the energy deposit occurs is tracked at each level:
@@ 5 The regularisation approach
This method does not aim at explicitly solving the initial equation, but rather to obtain a well-posed problem, similar to the starting ill-posed one, and solve it instead. Therefore, the solution obtained is an approximation. An example of such an approach is Tikhonov regularisation () [14]:
where the second part of the expression is known as a regularisation term. In this term, is a Tikhonov matrix and is the regularisation parameter, respectively. In other words, the expression (assuming that is a square matrix and invertible) does not satisfy Equation 3. The level of deviation is controlled by the regularisation parameter . This is the price to pay to obtain a unique solution. In Equation 8 the regularisation term is quadratic; its choice strongly depends on the nature of the initial problem. It is known [15, 16] that for sparse problems the so-called approach can be employed: it uses the absolute values of the target value (energies) as the regularisation term. The low multiplicity of cubes per plane makes the SoLid reconstruction a canonical sparse problem. Therefore, the loss function reads:
@@
where is the step size, sgn is a sign function and is a shrinkage operator. Recently, an upgraded version of the ISTA algorithm was developed. It has an improved complexity result of O(1/k2), which is achieved using the specific linear combination of the two previous iteration values inside the shrinkage operator. It is called FISTA [7] (for Fast ISTA) and was implemented in Saffron2 and tested.
The method was tested with a sample of 3105 IBD events. The convergence development is presented in Figure 4. With either of the calculation methods, convergence is improving with the number of iterations and reaching a plateau. The high value of data fidelity is explained by the fact that it is the sum of all events generated. The number of iterations was arbitrarily selected as thirty, and the physics estimators were obtained after this condition. They are summarised in Table 1 for several energy thresholds. As expected, the ghost-cube rate decreases drastically with larger energy cut values. However, a compromise has to be made in the choice of the energy threshold, since the reconstruction efficiency of the IBD annihilation gammas is decreasing accordingly. Dedicated optimisation must be performed in the analysis workflow, which is beyond the scope of this paper. Together with the reduction of the multiplicity of ghosts, the reconstruction efficiency improves with a larger threshold. This is another illustration that the algorithm copes much more easily with high-energy cubes.
@@ 6 The Bayesian approach
where denotes proportionality. In this approach, is the solution of Equation 3. The ML-EM method has the following set of features: accurate modelling of the Poisson noise distribution; the solution is constrained to be positive without regularisation; and unitarity of the sum of the projections is preserved (the total number of scintillation photons is conserved). The measured projections in SoLid obey the Poisson law. Therefore, the selection of the Maximum-Likelihood Expectation-Minimisation (ML-EM) method is one of the most adapted. It is widely used in medical imaging and lately in particle physics, e.g. in the NEXT experiment [8].
The algorithm iteratively generates the projected values based on the current approximation of the solution. Iterations are stopped at the desired precision. An initial guess of the answer is required at the beginning of the process. The choice of initial guess is critical because of the potential non-uniqueness of the solution; it is desirable to find the globally optimal solution. Two methods have been attempted to choose a starting point. First, a democratic approach that supposes an equal share of the total energy between the cubes. Second, the FISTA result described above is used. It has been shown [17] that the choice of the latter significantly improves, in general, the performance of ML-EM for sparse problems, and this is also observed in this work. Equation 12 defines the master formula of the ML-EM algorithm. It is obtained by maximising the probability of obtaining a measurement p.
@@ 7.1 Energy resolution
where is the antineutrino energy from the Geant4 simulation and is the energy of the most energetic cube per ROSim event. Events in the sample defined above are split into 25 bins with identical statistics. For each bin, the distribution is built and fit with a Crystal Ball function [18]. The fit result provides a value of , which is used to compute . In this case, is the median energy for each individual bin. The energy resolution studies for the same type of scintillator (EJ-200) are presented in [19]. It has been chosen to adopt the same opportunistic 3-parameter function to fit the data:

### v2

@@ Abstract
The SoLid experiment is a very-short-baseline experiment aimed at searching for nuclear reactor-produced active to sterile antineutrino oscillations. The detection principle is based on the pairing of two types of solid scintillators: polyvinyl toluene and LiF:ZnS(Ag), which is a new technology used in this field of Physics. In addition to good neutron-gamma discrimination, this setup allows the detector to be highly segmented (the basic detection unit is a 5 cm side cube). High segmentation provides numerous advantages, including the precise location of Inverse Beta Decay (IBD) products, the derivation of the considerate antineutrino energy estimator, and a powerful background reduction tool based on the topological signature of the signal. Finally, the system is read out by a network of wavelength-shifting fibres coupled to a photodetector (MPPC). This paper describes the design of the reconstruction algorithm that allows maximum use of the granularity of the detector. The goal of the algorithm is to convert the output of the optical-fibre readout to the list of the detection units from which it originated. This paper provides a performance comparison for three methods and concludes with a choice of the baseline approach for the experiment.
@@
0000, 1111MSC:
0000, 1111 2023 by SoLid collaboration is licensed under CC BY 4.0[inst1] organization=Université Clermont Auvergne, CNRS/IN2P3, LPCA, city=Clermont-Ferrand, country=France
@@ 1 Introduction
The SoLid experiment is located in the vicinity of the BR2 research reactor at the SCK CEN site in Mol, Belgium. The experiment aims to make a precise measurement of the reactor antineutrino flux at a very short baseline (6.7 - 9.2 m), as well as its energy spectrum, to study the “5 MeV bump” DoubleChooz:2019qbj . The first measurement allows, in principle, information on the so-called reactor anomaly RAA Mention_2011 and the Gallium anomaly Barinov_2022 to be provided. In particular, it is possible to test the 3+1 model Abazajian:2012ys , which assumes the existence of an additional light-sterile neutrinos state. The SoLid experiment was designed to probe the best-fit region of the oscillation parameters with sin() 0.1 and m = 1 eV.
@@
The description of the design of the SoLid detector is beyond the scope of this article. The interested reader can consult the detailed discussion reported in SoLid:2020cen . The following paragraph contains the executive summary required for the description of the reconstruction algorithm.
@@
The basic detection unit of the detector is a 5 cm side polyvinyl toluene (PVT) cube. The PVT is a plastic scintillator that acts as a proton-enriched target for the antineutrino and detects the light generated by the IBD positron and consequent annihilation gamma. The detection units are combined into planes of 1616 units each. Each cube is individually wrapped in Tyvek to prevent scintillation light from escaping. Furthermore, each plane is surrounded by two square Tyvek sheets, which optically decouple the planes by preventing light from passing between them even further (shown in Figure 1). The latter is a very important feature of the calibration algorithm. Ten planes make up a module, and there are five modules in total. The scintillation light from the PVT cube is collected by two vertical and two horizontal wavelength-shifting (WLS) fibres that pass through each detection cell. One side of each WLS fibre is covered with a Mylar foil that acts as a mirror to reflect the incoming light towards the other side. The second side is coupled with the multi-pixel photon counters (MPPCs) that read out the light. The digitised version of the readout received from the detector is the starting point for any further analysis. Therefore, signal candidates must be defined from this input. The most convenient and useful way to represent the data is to transform the MPPC response back into the list of detection cubes involved in the event. This type of problem, i.e. image reconstruction, has been investigated in medical physics, for example, in med_image_reco .
@@
Several algorithms have been applied for the reconstruction of SoLid data. Two of them are already well-known and widely used. The first is a regularisation approach called the Fast Iterative Shrinkage Threshold Algorithm or FISTA, which is an improved version of ISTA fista . The second is the Bayesian Maximum-Likelihood Expectation-Minimisation (ML-EM) approach, which is used e.g. in the NEXT experiment next . Finally, the last method is custom-made and builds on the ML-EM algorithm by augmenting it with information from the underlying physics processes at work. All three of them will be presented in this article, and their performances will be compared.
@@
One last important thing to mention is that all the results presented in this paper are obtained solely with Monte Carlo (MC) simulations. The production of simulation files in the SoLid experiment is based on several components. First, a detailed simulation of the core of the BR2 reactor is produced. It couples the reference ILL spectrum, fission rate predictions extracted from the MCNPX 3D model of the reactor core combined with the CINDER90 evolution code, and the MURE code to track the burn-up of the fissile products. Second, a detailed Geant4 Geant4:2002zbu simulation of the geometry of the SoLid detector is performed. Third, the SoLid Oscillation (SoLO) framework, which works as an antineutrino generator and mediator between the two simulations introduced above. For each signal event, SoLO peaks a position inside the reactor core according to the given fission map, an antineutrino energy from the predicted spectrum, and the interaction point inside the detector. By construction, this approach takes into account the geometrical acceptance of the detector. Thus, the computational cost is significantly reduced compared to isotropic generation. To reduce it even further, SoLO directly transfers the kinematic information of the positron and neutron to the detector simulation. The interested reader can consult the detailed description of the SoLO framework reported in Ianthe . The final piece that handles the output of the Geant4 simulation is the so-called ReadOut Simulation (ROSim). The goal of ROSim is to reproduce the response from the detector. Practically, it transforms the energy deposits into the number of scintillation photons that are read out by the correspondent MPPCs. The detailed tuning of the ROSim parameters is reported in maja .
@@ 2 Signal definition
The SoLid detector uses the IBD process to detect antineutrino. This justifies the use of two scintillators: LiF:ZnS(Ag) screens to capture the neutron and PVT as a proton-enriched target for the antineutrino, hence recording the positron scintillation. For the former, once the neutron is thermalised through elastic collision, it is captured and produces tritium and particles, detected in the ZnS(Ag) crystal, which is integrated in the neutron detection screens. The latter detects in addition to the positron scintillation the subsequent positron annihilation gammas. The antineutrino energy is defined from the IBD process as follows:
@@ 3 The CCube algorithm
Let us mathematically define the problem. By postulating a as the projector from cube j to MPPC i, the readout value is calculated as follows:
@@
Equation 3 is a discrete linear inverse problem. The problem is well-posed if it meets certain requirements: existence, uniqueness, and stability of the solution with respect to the small shifts. In the current case, the uniqueness condition is violated, as illustrated in Figure 3. For simplicity, each cube communicates there only through the fibres that cross it: the projectors a of the cubes crossed by the given fibre i are set to 1, and to 0 for all other pairs of i,j. This indicates that only two energy deposits in the same plane are enough to allow the readout to be fulfilled in several ways. In particular, false cubes may be created at the intersection of the fibres of the true ones. In addition to the multiplicity of events, false cubes also introduce an energy bias to the initial energy deposits. Thus, the reconstruction problem is ill-posed and shall be solved either by regularisation or Bayesian methods. Before discussing the methods themselves, let us define the set of estimators that are used to assess the performance of different methods.
@@ 4 Reconstruction procedure estimators
The estimators used to evaluate the performance of the various methods are divided into two categories. The first is related to the convergence of the algorithms only. The estimators in this group calculate the deviation (also indicated as data fidelity) of the solution derived (E) from the actual measurements (p). The least squares norm and the Kullback-Leiber divergence 10.1214/aoms/1177729694 were selected to obtain it. Their mathematical definition reads as follows:
@@
The second group of estimators involves physical performance. Geant4 (G4) simulation serves as truth-level information, while the output of the SoLid Analysis Framework (Saffron2 peste2019 or simply S2) with the reconstruction approach at work serves as a derived solution. The comparison relies on the list of cubes per event, their coordinates (positions in the plane) and associated energy contributions. To begin with, the total number of cubes in which the energy deposit occurs is tracked at each level:
@@ 5 The regularisation approach
This method does not aim at explicitly solving the initial equation, but rather to obtain a well-posed problem, similar to the starting ill-posed one, and solve it instead. Therefore, the solution obtained is an approximation. An example of such an approach is Tikhonov regularisation () tikhonov1977solutions :
where the second part of the expression is known as a regularisation term. In this term, is a Tikhonov matrix and is the regularisation parameter, respectively. In other words, the expression (assuming that is a square matrix and invertible) does not satisfy Equation 3. The level of deviation is controlled by the regularisation parameter . This is the price to pay to obtain a unique solution. In Equation 8 the regularisation term is quadratic; its choice strongly depends on the nature of the initial problem. It is known 4407762 ; 1217267 that for sparse problems the so-called approach can be employed: it uses the absolute values of the target value (energies) as the regularisation term. The low multiplicity of cubes per plane makes the SoLid reconstruction a canonical sparse problem. Therefore, the loss function reads:
@@
where is the step size, sgn is a sign function and is a shrinkage operator. Recently, an upgraded version of the ISTA algorithm was developed. It has an improved complexity result of O(1/k), which is achieved using the specific linear combination of the two previous iteration values inside the shrinkage operator. It is called FISTA fista (for Fast ISTA) and was implemented in Saffron2 and tested.
The method was tested with a sample of 310 IBD events. The convergence development is presented in Figure 4. With either of the calculation methods, convergence is improving with the number of iterations and reaching a plateau. The high value of data fidelity is explained by the fact that it is the sum of all events generated. The number of iterations was arbitrarily selected as thirty, and the physics estimators were obtained after this condition. They are summarised in Table 1 for several energy thresholds. As expected, the ghost-cube rate decreases drastically with larger energy cut values. However, a compromise has to be made in the choice of the energy threshold, since the reconstruction efficiency of the IBD annihilation gammas is decreasing accordingly. Dedicated optimisation must be performed in the analysis workflow, which is beyond the scope of this paper. Together with the reduction of the multiplicity of ghosts, the reconstruction efficiency improves with a larger threshold. This is another illustration that the algorithm copes much more easily with high-energy cubes.
@@ 6 The Bayesian approach
where denotes proportionality. In this approach, is the solution of Equation 3. The ML-EM method has the following set of features: accurate modelling of the Poisson noise distribution; the solution is constrained to be positive without regularisation; and unitarity of the sum of the projections is preserved (the total number of scintillation photons is conserved). The measured projections in SoLid obey the Poisson law. Therefore, the selection of the Maximum-Likelihood Expectation-Minimisation (ML-EM) method is one of the most adapted. It is widely used in medical imaging and lately in particle physics, e.g. in the NEXT experiment next .
The algorithm iteratively generates the projected values based on the current approximation of the solution. Iterations are stopped at the desired precision. An initial guess of the answer is required at the beginning of the process. The choice of initial guess is critical because of the potential non-uniqueness of the solution; it is desirable to find the globally optimal solution. Two methods have been attempted to choose a starting point. First, a democratic approach that supposes an equal share of the total energy between the cubes. Second, the FISTA result described above is used. It has been shown Zhu:18 that the choice of the latter significantly improves, in general, the performance of ML-EM for sparse problems, and this is also observed in this work. Equation 12 defines the master formula of the ML-EM algorithm. It is obtained by maximising the probability of obtaining a measurement p.
@@ 7.1 Energy resolution
where is the antineutrino energy from the Geant4 simulation and is the energy of the most energetic cube per ROSim event. Events in the sample defined above are split into 25 bins with identical statistics. For each bin, the distribution is built and fit with a Crystal Ball function crystalball . The fit result provides a value of , which is used to compute . In this case, is the median energy for each individual bin. The energy resolution studies for the same type of scintillator (EJ-200) are presented in ej200resolution . It has been chosen to adopt the same opportunistic 3-parameter function to fit the data:
@@ 8 Conclusion
9 Acknowledgements
This work was supported by the following funding agencies: Agence Nationale de la Recherche grant ANR-16CE31001803, Institut Carnot Mines, CNRS/IN2P3, and Region Pays de Loire, France; FWO-Vlaanderen and the Vlaamse Herculesstichting, Belgium; The UK groups acknowledge the support of the Science & Technology Facilities Council (STFC), United Kingdom; We are grateful for the early support given by the subdepartment of Particle Physics at Oxford and High Energy Physics at Imperial College London. We also thank our colleagues, the administrative and technical staff of the SCK CEN, for their invaluable support for this project. Individuals have received support from FWO-Vlaanderen and the Belgian Federal Science Policy Office (BelSpo) under the IUAP network programme; the STFC Rutherford Fellowship programme and the European Research Council under the European Union Horizon 2020 Programme (H2020-CoG)/ERC Grant Agreement n. 682474; Merton College Oxford.